OpenStudy (anonymous):

The length of a rectangle is 7 units and its width is 5 units. What is the approximate length of the diagonal of the rectangle? 4.9 units 6 units 8.6 units 9 units

OpenStudy (michele_laino):

we have to apply the theorem of Pitagora, namely: \[d = \sqrt {{L^2} + {W^2}} = \sqrt {{7^2} + {5^2}} = ...\] where d is the diagonal

OpenStudy (michele_laino):



OpenStudy (anonymous):

are u sure ? i havent heard of that in my lesson haha

OpenStudy (michele_laino):

yes! I'm sure!

OpenStudy (michele_laino):

what is: \[\Large \sqrt {{7^2} + {5^2}} = \sqrt {49 + 25} = ...\]

OpenStudy (anonymous):

74 ?

OpenStudy (michele_laino):

then you have to compute the square root of 74, so what do you get?

OpenStudy (anonymous):

8.6 ?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

OH LOL XD

OpenStudy (anonymous):

thank you <3

OpenStudy (michele_laino):

thank you! :)